OpenStudy (anonymous):

find the indicated one-sided limit. if the limiting value is inifinte indicate whther is is +∞or− , lim x→4 (3x−9)

OpenStudy (anonymous):

hi. i have a question. is the limit being taken from the left or right?

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

well, this is a linear function or a first degree polynomial.

OpenStudy (anonymous):

i'm not sure

OpenStudy (anonymous):

so tell me what the problem is asking, if you would be so kind.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 4}(3x^2-9) \]

OpenStudy (anonymous):

can you tell me in words, what it is asking?

OpenStudy (anonymous):

find the indicated one-sided limit. if the limiting value is inifinte, indicate whether it is \[+\infty or -\]

OpenStudy (anonymous):

-infity

OpenStudy (anonymous):

ok. so they want the limit of 3x^2-9 as x approaches 4. I see the 4, but is it a left- or right hand + limit. My glasses may not be picking up the detail.

OpenStudy (anonymous):

right hand

OpenStudy (anonymous):

ok. perfect. now, what do we know about 3x^2-9? what kind of a function is it?

OpenStudy (anonymous):

what is the graph of that function?

OpenStudy (anonymous):

i'm not sure...i'm a little confused on the whole topic of limits

OpenStudy (anonymous):

it is ok. it is a second degree polynomial. i am stating it this way because there is a limit theorem that states that all polynomial functions are continuous and all polynomial functions have equal left and right hand limits at every point in their domain.

OpenStudy (anonymous):

this function would graph out to be a parabola. Furthermore, unless special domains are part of the problem, you simply substitute the x value into the expression to get the limit.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

incidentally, polynomial functions are any function that can be written in the form: y=ax^n+bx^(n-1)+...+c

OpenStudy (anonymous):

so, 5x^6+3x^5+4=y is a sixth degree polynomial function. notice the degree and the highest power are the same.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok. the only reason i bring this up is because they tend to throw a lot of these polynomial functions at you in calc classes. now back to the problem at hand. did you substitute 4 for x in the expression?

OpenStudy (anonymous):

if so, what did you get for a value?

OpenStudy (anonymous):

39

OpenStudy (anonymous):

that would be the correct answer. nicely done.

OpenStudy (anonymous):

ok..thanks...its all starting to make sense now

OpenStudy (anonymous):

good. do you have any other questions? I am not paid by the hour lol

OpenStudy (anonymous):

lol. no not tonight

OpenStudy (anonymous):

well, then, i hope you have a lovely evening. take care and good nite

OpenStudy (anonymous):

thanks and the same to you

OpenStudy (anonymous):

:))

OpenStudy (anonymous):

what happens if the limit is being taken from the left?